OpenStudy (anonymous):

The path for a road race has the shape of a right triangle. For each lap, runners head due west from the starting point and run 1 mile. Then, they turn due north and run 2 miles, then return to the starting point. If the race includes 5 full laps, what is the total length of the road race? Round your answer to the nearest hundredth. A. 23.66 B. 26.18 C. 29.14 D. 30.81

OpenStudy (irishboy123):

from Pythagoreas: long leg = sqrt(1+4) = 2.24 race = 5*(2.24+1+2) = 26.18 = B

OpenStudy (anonymous):

thank you